Yahaira:

​Together, teammates Pedro and Ricky got 2682 base hits last season. Pedro had 276 more hits than Ricky. How many hits did each player​ have?

cathecat88:

Ricky has 1204 hits and Pedro has 1481 hits

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cathecat88 Ricky has 1204 hits and Pedro has 1481 hits \(\color{#0cbb34}{\text{End of Quote}}\) U cant give direct answers u need to delete that and explain how you get the answer

Yahaira:

It was wrong :/

Yahaira:

Thanks tho!, stop arguing

CripQUEZZ:

Ok so subtract 2682 by 276 2682 - 276 = 2406 Then to get half of it divide by two 2406/2 = 1203 So Ricky had 1203 And add 276 to 1203 1203 + 276 = 1479 So Pedro had 1479 and Ricky had 1203

Yahaira:

Thank you so much! I get it know!

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yahaira Thank you so much! I get it know! \(\color{#0cbb34}{\text{End of Quote}}\) Ok yw I was tryna help you get it right but they gave the wrong answer first

Yahaira:

its ok :) I'll understand the next one

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Yahaira its ok :) I'll understand the next one \(\color{#0cbb34}{\text{End of Quote}}\) Ok

CripQUEZZ:

If you need anymore help jus dm me and I can or @ me in a post